OpenStudy (jamesj):

For Tania08: At the beginning of the year, Quasar Company's liabilities equal $51,000. During the year, assets increase by $60,000, and at year-end assets equal $190,000. Liabilities decrease $12,000 during the year. What are the beginning and ending amounts of equity?

OpenStudy (anonymous):

I'm here now.

OpenStudy (jamesj):

Again, start with Assets = Liabilities + Equity BOY* _________ $51,000 _________ Change $60,000 ($12,000) _________ EOY** $190,000 _________ _________ *BOY = Beginning of Year **EOY = End of Year Now, fill in the gaps in this table.

OpenStudy (anonymous):

I'm lost. :/

OpenStudy (jamesj):

Assets = Liabilities + Equity. Agreed on that?

OpenStudy (anonymous):

Yes, I get that part.

OpenStudy (jamesj):

Good. Next, you're given four pieces of information: - liabilities at beginning of year: $51,000 - change in assets during year: $60,000 - assets at end of year: $190,000 - change in liabilities during year: ($12,000) .... where as usual ( ) around a number means negative So what I've done is create a table to show that information. That's what this is: Assets = Liabilities + Equity BOY* _________ $51,000 _________ Change $60,000 ($12,000) _________ EOY** $190,000 _________ _________

OpenStudy (jamesj):

talk to me ... tell me what's working and what's not

OpenStudy (jamesj):

talk to me or I'm out of here.

OpenStudy (anonymous):

Ok umm

OpenStudy (jamesj):

So given this information: Assets = Liabilities + Equity BOY* _________ $51,000 _________ Change $60,000 ($12,000) _________ EOY** $190,000 _________ _________ What are the assets at the beginning of the year, BOY?

OpenStudy (jamesj):

You know what the assets are at the end of the year (EOY), and you know how much they changed. Hence what were they at the BOY?

OpenStudy (jamesj):

Let X = assets at the beginning of the year. X + $60,000 = $190,000 Hence what is X?

OpenStudy (jamesj):

X must be $130,000. If you end the year with $190,000 of assets, and during the year the assets changed by $60,000, it must be that we started the year with $130,000 of assets. Hence our table now looks like this: Assets = Liabilities + Equity BOY* $130,000 $51,000 _________ Change $60,000 ($12,000) _________ EOY** $190,000 _________ _________ Now, given that, you can calculate the equity at the beginning of the year, one of the two things the question originally asked you for. In fact, you should really have no trouble filling in every blank in this table.

OpenStudy (anonymous):

Can you just tell me what I have to add and subtract?

OpenStudy (jamesj):

Well, Assets = Liabilities + Equity Hence what is Equity on this line: Assets = Liabilities + Equity BOY* $130,000 $51,000 _________

OpenStudy (anonymous):

I need two answers though.

OpenStudy (jamesj):

I know ... and this is how you get there.

OpenStudy (jamesj):

Figure out how to complete the table. Think it through and be logical. Complete each row in turn. It's really not hard to figure out what to subtract and what to add.

OpenStudy (jamesj):

For example, with the first row, we have Assets = Liabilities + Equity BOY* $130,000 = $51,000 + ______ What value of Equity makes sense here?

OpenStudy (anonymous):

Is the answer of equity $79.000?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

Sorry, I´ve answered just the inicital equity. It follows the right answers: initial equity = 79.000,00 and the final equity = 127.000,00. You have the following data: initial liabiliaties (IL), the asset´s variation (VA), the final asset (FA) and the liabilities variation (VL). So, Initial asset (IA)= FA - VA = 190.000-60.000 = 130.000. Inital equity = IA- IL = 130.000-51.000 = 79.000. Using the same reasoning, you will get the final equity, because the final liabilities (FL) = IL+VL = 51000+12000 = 63.000. The final equity = FA-FL = 190.0000-63.000 = 127.000. Did you realize that the equity´s variation is equal to asset´s variation minus liabilities variation, i.e, 48.000? You got?

OpenStudy (anonymous):

SO its 127000?

OpenStudy (anonymous):

There was two questions: the value of equity at the beggining and at the end. At the beggining the answer is 79.000, and at the end the answer is 127.000.

OpenStudy (anonymous):

127000 is wrong. =/

OpenStudy (anonymous):

Ok. The problem was that the liabilities " decreased"...